# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v9

@@ Abstract
[Objective] This study addresses key challenges in machine learning, namely the absence of a unified formal theoretical framework and the lack of foundational theories for model interpretability and ethical safety.
[Methods] We first construct a formal information model, explicitly defining the ontological states and carrier mappings of typical machine learning stages using sets of well-formed formulas. By introducing learnable and processable predicates, as well as learning and processing functions, we analyze the causal chain logic and constraint laws governing machine learning processes.
[Results] We establish the Machine Learning Theory Meta-Framework (MLT-MF), on which we further propose universal definitions for model interpretability and ethical safety. We prove and validate three key theorems: the relationship between model interpretability and information existence, ethical safety assurance, and the upper bound estimation of total variation distance (TVD).
[Limitations] The current framework assumes ideal, noise-free information enabling mappings and focuses primarily on model learning and processing logic in static scenarios. It does not yet address information fusion and conflict resolution across ontological spaces in multimodal or multi-agent systems.
[Conclusions] This work overcomes the limitations of fragmented research and provides a unified theoretical foundation for systematically addressing critical issues in contemporary machine learning.
Keywords: machine learning; objective information theory; formal logic systems; interpretability; ethical and safety assurance; generalization error
@@ 1 Introduction
Although artificial intelligence is currently the most dynamic field within information science and even the broader domain of science and technology, existing research still faces several significant shortcomings:
First, there is a lack of a formalized framework that covers the entire lifecycle of machine learning Rob2021 ; Marius2023 ; Meng2023 ; Alquier2024 ; Vilalta2002 . Current theoretical studies in machine learning tend to focus on isolated components: optimization theory emphasizes training dynamics, generalization analysis concerns test performance, and strategies for initialization and deployment are often treated as independent stages. This fragmentation makes it impossible to characterize the coherent evolution of knowledge from prior injection, through data-driven updates, to task generalization. For example, the PAC-Bayes framework connects priors with generalization error but neglects the temporal causality of the training process; meta-learning involves multi-stage adaptation but fails to establish a symbolic model for state transitions. The absence of lifecycle modeling makes systematic analysis of the core question——“how parameters carry task knowledge from initialization through training”—difficult for model designers.
Second, there is a lack of explicit expression for the causal chain of information flow. Traditional approaches often treat the transformation from data to model as a black box, reducing the parameter update driven by loss function gradients to a purely mathematical optimization process and overlooking the causal dependencies among prior knowledge, data distribution, and model architecture. While the field of causal inference attempts to characterize causal relationships among variables, its application in machine learning is mostly limited to feature-label association analysis and fails to construct a complete causal chain from “prior constraints → initialization → gradient propagation → generalization capability.” For instance, adversarial training enhances robustness by maximizing loss, but theoretical analyses rarely explicitly link initial parameter constraints with the causal mechanisms underlying final adversarial robustness, making strategy design reliant on empirical trial-and-error.
Third, the mathematical foundations of interpretability tools are weak Ribeiro2016 ; Lundberg2017 ; Venkata2025 ; Lee2023 ; Alvarez2018 . Most current interpretability methods use local approximation or game-theoretic approaches to assign feature importance, which, while practically useful, lack rigorous mathematical axiomatization. For example, SHAP values compute marginal contributions in cooperative games but do not formalize their mapping to the information carrier states of models, potentially violating causality in interpretations. In addition, heuristic algorithms that infer data-feature dependencies from parameters lack mathematical guarantees for “information recoverability.” This theoretical weakness makes interpretability results susceptible to adversarial attacks and limits their applicability in high-stakes scenarios such as medicine or law.
Information science, as a core discipline connecting data, knowledge, and systems, offers a new perspective for the theoretical integration of machine learningShwartz2017 ; Xu2022 . Objective Information Theory (OIT) establishes a sextuple model of information through axiomatic methods, representing the intrinsic meaning of information via ontological states, reflecting its objective expression through carrier states, and explaining information realization through enabling mappings that possess both mathematical surjectivity and physical causality. This provides a universally applicable foundational model for understanding information in the objective world, especially within computer and information systemsXu2014 ; xu2015 ; Jianfeng2021 ; Jianfeng2023 ; Xu2024 ; Xu2025 .
On this basis, we rigorously define higher-order logic systems using recursion, and establish formal state expressions through interpretations of sets of well-formed formulas (WFFs), completing the last theoretical gap in OIT regarding the mathematical essence of state expressions. We then thoroughly apply the theory and methods of OIT to unify the core concepts throughout the machine learning lifecycle—prior knowledge encoding, data-driven training, and model generalization—as formal state expressions; we establish a mathematical logic system covering the entire ML process. Using this, we analyze and prove the causal-chain logic and constraint rules for typical stages in machine learning, including the formal modeling and realization of training information, the formal modeling and realization of model learning, and the model’s information processing and learning optimization, thus forming a Meta-Framework for Machine Learning Theory (MLT-MF). Finally, using MLT-MF, we propose universal definitions of model interpretability and ethical safety, prove and validate—via a simple feedforward neural network—three important theorems: model interpretability and information existence, ethical safety assurance, and an upper bound estimate on total variation distance (TVD). These results provide a powerful tool, based on a unified formal logic framework, for both theoretical analysis and engineering practice in machine learning.
2 Postulates and Formal Definition of Information
@@
2.1 Postulates of Information
Although humanity has entered the information age, the theoretical foundations of information science have long been dispersed across disciplines such as information theory, cybernetics, signal processing, statistical learning, and computational theory, each developing independently. There remains a lack of consensus on the concept of information, as well as a unified axiomatic basis and theoretical system. This is also a major reason why artificial intelligence often lacks interpretability and predictability.
To address this, we first propose the concept of an enabling mapping.

### v10

@@ Abstract
This paper addresses the current lack of a unified formal framework in machine learning theory, as well as the absence of robust theoretical foundations for interpretability and ethical safety assurance. We first construct a formal information model, employing sets of well-formed formulas (WFFs) to explicitly define the ontological states and carrier mappings for the core components of machine learning. By introducing learnable and processable predicates, as well as learning and processing functions, we analyze the logical inference and constraint rules underlying causal chains in models, thereby establishing the Machine Learning Theory Meta-Framework (MLT-MF). Building upon this framework, we propose universal definitions for model interpretability and ethical safety, and rigorously prove and validate four key theorems: the equivalence between model interpretability and information existence, the constructive formulation of ethical safety assurance and two types of total variation distance (TVD) upper bounds. This work overcomes the limitations of previous fragmented approaches, providing a unified theoretical foundation from an information science perspective to systematically address the critical challenges currently facing machine learning.
Keywords: machine learning, objective information theory (OIT), formal logic systems, interpretability, ethical safety assurance, estimation of error bound
@@ 1 Introduction
Although artificial intelligence (AI), exemplified by machine learning (ML), is among the most dynamic areas in information science and technology, several critical shortcomings persist in the current research landscape. First, there is a lack of a unified logical framework that spans the entire lifecycle of machine learning. Most existing ML theory focuses on isolated components: optimization theory emphasizes training dynamics, generalization analysis targets test performance, while initialization strategies and deployment are often considered independently. This fragmentation precludes the coherent characterization of knowledge evolution from prior injection, data-driven updates, to task-level generalization (Vapnik, 1999; Alquier, 2024; Schölkopf et al., 2021; Garcez and Lamb, 2023; Marra et al., 2024).
Second, the unpredictability of model performance has led to rapidly escalating research and development costs. As model sizes continue to grow, the demands for data, computation, energy, time, and human resources have multiplied. In the absence of explicit causal semantics, structural analysis, and effective metrics, it is often challenging to analyze or predict attainable performance in each stage of the pipeline based on error upper bounds. Incomplete information coverage, suboptimal model selection, inaccurate measurement, and quality deficiencies frequently result in failed development efforts and significant resource losses (Achiam et al., 2023; Touvron et al., 2023; Chowdhery et al., 2023; Gunasekar et al., 2023).
Third, there are profound concerns regarding ethical safety assurance. As AI capabilities advance, so too do apprehensions about their ethical implications. However, there is still no consensus—either mathematical or theoretical—on the scientific definition, substantive content, forms of expression, or criteria for ethical safety in AI. Most research remains focused on ethical principles and regulatory mechanisms, while the proliferation of competing guidelines leaves practitioners uncertain about which standards to follow (Russell et al., 2015; Burr and Leslie, 2023; Tilala et al., 2024; Hawkins et al., 2021; Pant et al., 2024).
Fourth, the theoretical foundations for interpretability remain weak. The lack of a unified semantic framework for causal inference and constraint logic means that traditional ML methods treat the transformation from data to model as a black box. Although interpretability is widely recognized as a critical issue and diverse approaches have been proposed to address it, there is as yet no rigorous definition or universally applicable evaluative criteria (Lipton, 2018; Glanois et al., 2024; Zhang et al., 2024; Rudin, 2019; Miller, 2019).
As a central discipline bridging data, knowledge, and systems, information science has begun to offer new perspectives for the theoretical integration of machine learning (Shwartz-Ziv and Tishby, 2017; Xu et al., 2022). However, information science theory remains fragmented across information theory, control theory, signal processing, statistical learning, and computational theory, each developing largely in isolation and lacking a unified semantic and logical system. This fragmentation is a primary reason for the absence of a unified theoretical framework in contemporary AI. In response, Objective Information Theory (OIT) applies an axiomatic approach to construct a sextuple model of information, representing intrinsic meaning via ontological states, objective expression via carrier states, and revealing the dual-subject structure of information through the enabling mapping from the former to the latter. OIT thus provides a general and universal foundational model and metrics for understanding information in the objective world, and especially within computing and information systems (Xu et al., 2014; Xu, 2024). For example, (Xu et al., 2025) demonstrates that applying OIT’s general information metrics to pre-evaluate training data can significantly enhance model training efficiency.
This paper further defines the enabling mapping, clarifying its dual mathematical and physical-causal attributes. Through a recursively defined higher-order logical system, we formalize the expression of states via the interpretation of sets of well-formed formulas (WFFs), elucidating the mathematical essence of state representation in information definitions. This fills the final theoretical gap in OIT and offers a more powerful toolkit for applying OIT to diverse problems in information science. The introduction of noisy information and information chains further establishes the theoretical groundwork for analyzing key challenges across all stages of machine learning.
On this basis, we comprehensively apply OIT’s theories and methodologies—centered on the learnable and processable predicates, as well as learning and processing functions—to five canonical stages in machine learning: model construction, training information, learning and training, input of problems, information processing outputs and online learning. We establish a unified semantic formalization of state representation, and develop causal-chain logical inference and constraint rules covering the full ML lifecycle, culminating in the Meta-Framework for Machine Learning Theory (MLT-MF).
Applying MLT-MF, we formulate universal definitions for model interpretability and ethical safety, and prove theorems relating interpretability to information existence, as well as the theoretical underpinnings for ethical safety assurance. For model error, we establish total variation distance (TVD) upper bounds for two strategies—conservative prediction in regions not covered by training with noiseless input, and arbitrary prediction with noisy input—providing new guidelines for the estimation and optimization of machine learning performance. Finally, we illustrate and validate the practical utility of MLT-MF and the associated theorems with a simple feedforward neural network model, demonstrating that this research offers a novel direction and methodology for both theoretical analysis and engineering practice in machine learning, grounded in a unified formal logical framework.
2 Related Work
2.1 Conceptual Representations of Machine Learning
Generally, machine learning refers to methods that learn general rules from limited observational data and use these rules to predict unknown information. Numerous works have provided more detailed analyses of this concept. Representative views include:
In the context of statistical learning (Vapnik, 1999), the learning model is decomposed into three components: a set of independent random vectors $x$ that follow a certain but unknown probability distribution $P(x)$; a set of output vectors $y$ corresponding to input $x$, also following a certain but unknown conditional probability distribution $P\left(y\middle|x\right)$; and a collection of functions $f(x,\alpha),\alpha\in\Lambda$, that the learning machine can implement. Machine learning is thus defined as selecting, based on a training dataset composed of $l$ independent observations $\left(x_{1},y_{1}\right),\cdots,(x_{l},y_{l})$ drawn from the same distribution $P(x,y)=P(x)P\left(y\middle|x\right)$, a function $f(x,\alpha)$(with $\alpha\in\Lambda$) that minimizes the risk function $R(\alpha)=\int{L(y,f(x,\alpha))d}P(x,y)$, where $L(y,f(x,\alpha))$ is the loss function of $f(x,\alpha)$.
Another perspective highlights the central role of model parameters (Alquier, 2024), describing machine learning as the process of learning from an object-label set $(X,Y)$ governed by a probability distribution $P$, to produce a family of predictors $\left\{f_{\theta},\theta\in\Theta\right\}$ parameterized by $\theta\in\Theta$. The loss function $\mathcal{l(}f_{\theta}(X),Y)$ quantifies the performance of each predictor, and the expected risk is defined as $R(\theta)=\mathbb{E}_{(X,Y)\sim P}\mathcal{[l(}f_{\theta}(X),Y)]$. The ultimate goal is to find $\theta^{*}$ such that $R\left(\theta^{*}\right)=\inf_{\theta\in\Theta}R(\theta)$.
From a more advanced perspective, the focus shifts to causal relationships beyond data distribution features (Schölkopf et al., 2021). This view treats perceptual data acquired through measurement as entangled observations of unknown causal system states. It seeks to learn concrete representations that at least partially reveal these hidden causal structures by leveraging large-scale training, data augmentation, self-supervised learning, and robust fine-tuning using data from multiple simulated environments. This includes discovering causal models implicit in environmental data and performing causal inference to plan and act based on these models.
Some researchers emphasize the integration of neural and symbolic artificial intelligence (Garcez and Lamb, 2023), arguing that knowledge should be rooted in efficient data learning via message passing in neural networks and embedded in vector representations as efficient computational models. Symbols, in turn, should be acquired through querying and extracting knowledge from trained networks, providing a rich descriptive language at appropriate levels of abstraction—allowing finite means to achieve infinite ends—while enabling compositional and discrete reasoning at the symbolic level, thereby permitting extrapolation beyond data distributions. The combination of learning and reasoning should offer an important alternative for tackling compositional reasoning problems, reducing the number of effective combinations through learning and thus generating simpler symbolic descriptions within a neuro-symbolic cycle.
Another viewpoint proposes that structure learning through parameter learning is the most prominent paradigm among the statistical relational learning approaches within neuro-symbolic AI (Marra et al., 2024). The learning task involves estimating probabilities for a fixed set of clauses, with the advantage of removing combinatorial search from the learning process.
The above works collectively reflect, from different perspectives, the academic community’s comprehensive understanding of both past achievements and future directions in machine learning. However, none have established a formalized expression for the entire lifecycle implementation process of machine learning under a unified semantic framework. As a result, it remains consistently difficult to apply structured systems analysis methods to guide the implementation and optimization of models throughout research and design.
2.2 Model Interpretability
The interpretability of machine learning has attracted significant attention in recent years. As noted in (Lipton, 2018), few authors in the academic literature are able to precisely articulate what interpretability actually means, or to clearly explain how their proposed solutions achieve it. Despite the absence of a unified definition, an increasing number of publications continue to introduce algorithms purported to possess interpretability. This situation supports two possible conclusions: first, that there is a universally accepted understanding of interpretability that simply remains unwritten; or second, that the term itself is inherently ambiguous, leading to quasi-scientific claims regarding the interpretability of various models. In reality, the objective of interpretability may simply be to extract more valuable information from models.
Reference (Glanois et al., 2024) argues that, in the context of reinforcement learning, what is considered understandable and what constitutes a good explanation may differ entirely depending on whether the end-user, system designer, or policymaker is being considered. Based on a definition of comprehensibility encompassing simulatability, decomposability, and algorithmic transparency, the authors emphasize that interpretability in reinforcement learning relies on three pillars: interpretable inputs, interpretable models, and interpretable decisions. They further analyze the specific mechanisms for achieving interpretability in each of these aspects.
In (Zhang et al., 2024), the impact of different training mechanisms on two key factors—model smoothness and gradient similarity—is investigated from the perspective of agents, with the aim of deepening our understanding of the dynamics of model transferability and its relationship to various training methods. Building on adversarial training, two working hypotheses are proposed and validated: first, that the trade-off between model smoothness and gradient similarity largely determines the adversarial transferability of agent models; and second, that shifts in data distribution can undermine gradient similarity, thereby resulting in generally poorer substitutes. This approach—explaining changes in model performance by dissecting and adjusting key factors in the training process—should be regarded as an important direction in the study of machine learning interpretability.
It is evident, therefore, that although interpretability is a critical issue in ML, the lack of a unified semantic and formal framework for machine learning has precluded the establishment of rigorous definitions and universally applicable criteria for interpretability. Approaches to enhancing interpretability remain diverse, and there is still no universally accepted methodology informed by a common theoretical foundation.
2.3 Ethical Safety Assurance
Concerns regarding the ethical safety of AI have a long history. As noted in [10], the proliferation of intelligent and autonomous systems has given rise to significant legal and moral questions, the answers to which impact both producers and consumers of AI technologies. These issues span law, public policy, professional ethics, and philosophical morality, necessitating the expertise of computer scientists, legal scholars, political scientists, and ethicists. While mathematical tools such as formal logic, probability theory, and decision theory have provided important foundations for reasoning and decision-making, numerous challenges remain unresolved.
Building on traditional concepts of assurance, [11] offers a definition of ethical assurance for AI: it is a process that employs structured argumentation to provide auditable guarantees—specifically, that a set of normative claims regarding a system’s ethical properties are justified in light of the available evidence. Accordingly, the general structure of moral assurance is characterized as an iterative process of reflection, action, and adjustment, repeated throughout the design, development, and deployment lifecycle.
In the context of AI/ML-supported healthcare, ethical safety assurance faces more concrete challenges. Reference [12] provides a comprehensive exploration of the multifaceted ethical considerations in deploying AI in healthcare. The authors advocate for multiple strategies, including the implementation of rigorous data security protocols, prioritization of data minimization, increasing diversity in training datasets, use of data augmentation to address imbalance, application of algorithmic fairness techniques to detect and mitigate model bias, regular auditing of learning algorithms, enhancing the interpretability of algorithmic outputs, regulatory oversight based on standardized performance metrics and stringent testing protocols, and the augmentation—rather than replacement—of clinicians’ capabilities as means to achieve ethical safety.
[13] further investigates safety assurance guidelines for machine learning in specific domains such as autonomous systems, establishing a causal chain from ML safety to software safety and ultimately to system safety. The authors propose that ML safety requirements should be derived from the system-level safety requirements, and that these should be evaluated in light of the allocated safety needs, system and software architecture, and operational environment. This framework then serves as the basis for developing safety assurance specifications across all stages of the ML lifecycle, including data preparation, model training, model validation, and model deployment.
Reference [14] conducts an extensive study of the technology practitioners who have the greatest influence on AI ethical safety, proposing a classification framework that encompasses practitioner awareness, perceptions, needs, challenges, and approaches. The study highlights that, in the absence of a universally adopted set of ethical principles for AI, incorporating morality into AI is fundamentally challenged by the lack of clear moral knowledge and the ambiguity of ethical principles.
It is thus evident that, despite widespread concerns over the ethical safety of machine learning, there is as yet no mathematical or theoretical consensus regarding its scientific definition, substantive content, forms of manifestation, or criteria for assessment. Instead, ethical safety remains an open-ended concept, explored through various technical, engineering, and moral perspectives. Most research either focuses on specific ethical principles and consequences for AI, or on methods for enhancing ethical conduct. The proliferation of often conflicting AI ethical guidelines has, paradoxically, made it difficult for practitioners to determine which standards to follow. This underscores the necessity of a universally accepted definition of ethical safety, as such a definition can bridge the gap between ongoing academic discourse and industry practice, thereby providing a concrete foundation for strengthening ethical safety assurance in AI.
2.4 Model Error Estimation
Model performance is fundamentally reflected in the errors it may incur on specific tasks. Reference (Vapnik, 1999) provides key theorems on the consistency of empirical risk minimization (ERM); it characterizes the conditions for both two-sided and one-sided uniform convergence from empirical risk to true risk using VC entropy, VC dimension, and growth functions. Based on VC dimension, it further establishes distribution-independent risk upper bounds and sufficient conditions for fast convergence under both bounded and unbounded loss in finite-sample settings. By controlling the trade-off between VC dimension capacity and empirical risk through nested hypothesis families, it proves the strong consistency and convergence rates of structural risk minimization (SRM). These theoretical foundations have led to practical implementations such as support vector machines (SVM), margin maximization, and kernel methods, offering explanations for their generalization mechanisms.
Reference (Rudin, 2019) replaces global complexity with local empirical Rademacher averages, measuring complexity on subsets with small empirical error. It provides finite-sample error upper bounds via the fixed point of sub-root functions, thus achieving optimal or near-optimal convergence rates inversely proportional to the square root of the number of training samples. Under controlled variance conditions, it presents a unified framework for error estimation, including both distribution-dependent bounds determined by local Rademacher complexity and data-dependent bounds constructed from empirical local Rademacher averages. These results are extended to typical classification and regression tasks, providing computable upper bounds and iterative algorithms for quickly approximating the fixed point.
From an information-theoretic perspective, (Miller, 2019) models learning algorithms as random channels mapping input datasets $S$ to output hypothesis spaces $W$. For sub-Gaussian (including bounded) losses and uncountable hypothesis spaces, the generalization error upper bound is characterized via the mutual information $I(S;W)$ between input and output. Based on this, high-probability sample complexity and expected generalization error upper bounds are derived, and the concept of information stability $(\varepsilon,\mu)$-stability is introduced, unifying algorithmic consistency and overfitting control through information leakage measures. The paper further analyzes two classes of algorithms—Gibbs learning and noisy ERM—where $I(S;W)$ is explicitly controlled via mutual information regularization, and discusses methods for decomposing and additively controlling mutual information in complex learning algorithms through adaptive combinations.
Reference (Shwartz-Ziv and Tishby, 2017) provides a foundational theoretical framework for meta-reinforcement learning, focusing on in-distribution cross-task generalization and training convergence. It introduces a task distribution T and task-level MDPs, defines generalization error as the performance gap between training and unseen tasks under a meta-policy, and establishes generalization upper bounds determined by the number of training tasks and task complexity, using empirical process theory and concentration inequalities. The analysis further discusses the impact of algorithmic capacity and learning rates on error bounds. Within the non-convex SGD framework, sufficient conditions and rates for finite-step convergence of meta-learning loss are presented, along with guidelines for long-term behavior and hyperparameter selection based on linear and sub-linear types. The theoretical results are validated and the observability of convergence rates is demonstrated through controllable task generation and baseline meta-algorithms.
Reference (Xu et al., 2022) addresses the separation and convergence complexity of continuous and discrete errors in deterministic sampling of generative models. From both theoretical and numerical perspectives, it investigates the convergence properties of deterministic samplers based on probability flow ordinary differential equations (ODEs). Assuming access to score function estimates with $L_{2}$ accuracy, it provides upper bounds on TVD between target and generated distributions in continuous time. For the practical implementation using Runge-Kutta integrators, discrete-time error upper bounds are established. Numerical experiments on 128-dimensional problems are conducted to validate these theoretical error bounds.
Collectively, these works encompass both classical statistical learning theory and recent advances in reinforcement learning and generative models, reflecting a broad recognition of the importance of model error upper bounds in machine learning. However, these studies do not employ explicit causal semantic modeling, nor do they address the full lifecycle challenges of multi-stage information transmission, online learning, or multi-agent integration in ML. Additionally, they rarely consider the geometric scale and entropy of distribution shifts and uncovered domains, making it difficult to offer actionable engineering or data-sampling guidance for test domain coverage gaps and noisy input scenarios.
3 Preliminary Theory
@@
3.1 Definition of Information
Reference (Xu et al., 2014), based on four fundamental postulates, proposes that information is an enabling mapping from the state of an ontology to the state of a carrier, but does not further elaborate on the notions of “state” or “enabling mapping.” Appendix A establishes a higher-order formal system $\mathcal{L}$ and, drawing on general common sense, formulates basic postulates regarding states. It is then proven that any set of states can be represented as the interpretation of a set of WFFs in $\mathcal{L}$ (Xu, 2024), indicating that both ontological states and carrier states correspond to interpretations of sets of WFFs over specific domains of objects and times. Here, we proceed to further clarify the concept of enabling mapping.